OpenStudy (anonymous):

How do I find dx/dy by implicit differentiation and evaluate the derivative by the given point? x cos y = 1 ; (2, π/3)

zepdrix (zepdrix):

Since we'll be taking our derivative with respect to x, each time we differentiate a `non-x term`, we tack on a y'.\[\Large x \cos y = 1\] On the left side we'll apply the `product rule`:\[\Large \color{royalblue}{(x)'}\cos y+x \color{royalblue}{(\cos y)'}=\color{royalblue}{(1)'}\]

zepdrix (zepdrix):

So we need to take the derivative of the blue stuff.

zepdrix (zepdrix):

derivative of 1?

OpenStudy (anonymous):

The derivative of 1 is 0

OpenStudy (anonymous):

But how do I evaluate the derivative at the point (2, π/3)

zepdrix (zepdrix):

1 gives you 0? sounds right: \[\Large \color{royalblue}{(x)'}\cos y+x \color{royalblue}{(\cos y)'}=\color{orangered}{0}\] Understand how we calculated the left side?\[\Large \color{orangered}{(1)}\cos y+x \color{orangered}{(-\sin y)y'}=\color{orangered}{0}\]

zepdrix (zepdrix):

Solve for y':\[\Large y'\quad=\quad \frac{\cos x}{x \sin y}\]See how the right side involves some x and y stuff? We can think of y' as a function of both x and y. \[\Large y'(x,y)\quad=\quad \frac{\cos x}{x \sin y}\]And from here they want you to evaluate the derivative at x=2, y=pi/3.\[\Large y'\left(2,\;\frac{\pi}{3}\right)\quad=\quad?\]